40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 11 03:14:19 2013, Jenne, Update, LSC, Yarm held nicely on IR resonance with ALS, PRMI+arm attempt Yarm_onIRresonance_noPRMIyet_POYcalibrated.pdf
    Reply  Thu Jul 11 07:34:42 2013, manasa, Update, LSC, Yarm held nicely on IR resonance with ALS, PRMI+arm attempt 
Message ID: 8825     Entry time: Thu Jul 11 03:14:19 2013     Reply to this: 8826
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Yarm held nicely on IR resonance with ALS, PRMI+arm attempt 

[Annalisa, Jenne, Nic]

After having troubles with the Xarm earlier (maybe Manasa can write/say something about this?  Something about perhaps seeing the phase tracker jump, and cause it to lose lock?), we moved on to the Y arm. 

Annalisa locked the Yarm green, and closed the ALS loop.  I believe that earlier today, she tuned the gain such that we don't start getting gain peaking at a few hundred Hz.  We would like to get a script going, so that it's not so labor intensive to reclose the ALS loop after an MC lockloss....but that's a daytime task.

We then found the IR resonance, using only the Yarm ALS system.  After Manasa's work yesterday, the Yarm was very stable while locked with the ALS.  We took a power spectrum of POY11_I_ERR, which I have calibrated using the number in elog 6834 of 1.4e12 cts/m, or 7.14e-13 m/ct.  See the figure below.

After that, we changed the offsetter2 offset such that the arm was off resonance, but not so far off that we crossed any significant resonances (in particular, we wanted to not go as far as the 55MHz resonance). 

Then, I tried to lock the PRMI for a while, but the alignment wasn't very good.  We knew that the Yarm was well aligned, since our IR resonance was > 0.98, but it had been a while since we had aligned the X arm.  I tweaked the ITMX position to make the Michelson dark, and then tried acquiring PRMI lock.  At first, I tried with REFL165 I and Q, but with the non-ideal alignment and the offset in the 165 diode (LSC offsets was not run this evening), I wasn't catching any locks.  I then switched to AS55Q and REFL33I, but wasn't able to catch lock there either. 

The MC lost lock, which made us lose the ALS loop, but the ALS had been locked for more than 30 minutes, at least.  I tried locking the PRMI with the current alignment (after having misaligned ETMY), but was only able to get lock stretches of 1 second at maximum.

We are calling it a great success for the night, since we have confirmed that, at least for the Yarm, Manasa's beatbox work has improved things.  Also, we have a pretty solid plan for trying the PRMI+arm tomorrow, even though it didn't work out tonight.

Attachment 1: Yarm_onIRresonance_noPRMIyet_POYcalibrated.pdf  22 kB  | Hide | Hide all
Yarm_onIRresonance_noPRMIyet_POYcalibrated.pdf
ELOG V3.1.3-